OpenStudy (anonymous):

If sin (x) = 04.2 What's sin(pi+x) answer is -0.42 Please show working out, i dont get how they got that answer T_T

OpenStudy (mathteacher1729):

Do they want you to use the angle addition formula? It would be much easier to argue from looking at the graph of sin(x), though. :)

OpenStudy (anonymous):

It doesnt say anything. But the exercise im on is "Symmetry properties of circular functions"

OpenStudy (anonymous):

i don't think you have to get how it is done. it is a property

OpenStudy (anonymous):

Look at the pic. The functions that are positive in each quadrant are marked. If u start in first quadrant and add pi, u end up in third quadrant.

OpenStudy (anonymous):

The whole thing is 2pi of course (circle).

OpenStudy (mathteacher1729):

Go to http://www.wolframalpha.com and type in plot sin(x) and sin(x+pi) from x = 0 to 2pi You should see that they are mirror images of one another about the x-axis. In other words, whatever your y-coordinate is for sin(x) , the y-coordinate will be the same, but negative for sin(x+pi). :)

OpenStudy (anonymous):

you are looking directly across the unit circle, that is why.

OpenStudy (anonymous):

take a look at unit circle on last page of cheat sheet. pick any point. then traverse half way across the circle. it will be totally clear

OpenStudy (anonymous):

sine is a periodic function, meaning it repeats its y values after every interval. in the equation f(x)=sin (x+n), if n=2kpi (k is an integer), sinx=sin(x+n). if n=(2k+1)pi, sinx=-(sin(x+n))

OpenStudy (anonymous):

n=(2k)pi: even integer. n=(2k+1)pi: odd